OpenStudy (anonymous):

Show that (cos(x)) + sin(x))^2 = 1+ sin(2x) for every number x.

OpenStudy (anonymous):

\[ \begin{align} \big((cos(x)+sin(x)\big)^2&=1+sin(2x)\\ \big((cos(x)+sin(x)\big)\cdot\big((cos(x)+sin(x)\big)&=\\ cos^2(x)+2sin(x)cos(x)+sin^2(x)&=\\ 2sin(x)cos(x)+1&=1+sin(2x)\\ 2sin(x)cos(x)&=sin(2x) \end{align} \] This proof uses two fundamental trig identities: \[ sin^2(x)+cos^2(x)=1\\ 2sin(x)cos(x)=sin(2x) \]

OpenStudy (anonymous):

can you re-write this, I don't understand what you are writing. Sorry.

OpenStudy (anonymous):

this is (Cosx + sinx)^2 = 1+sin2x sin^2x + cos^2x =1 ( this is an identity ) open the bracket using (a+b)^2 = a^2 +2ab +b^2 Cos^2x +2CosxSinx + Sin^2x now we know that Cos^2x +Sin^2x=1 right? this means 2CosxSinx +1 Now there is another double angle identity which says that Sin2x=2CosxSinx So Sin2x+1= Right hand side as in the question and Hence its proved